OpenStudy (horsegal244):

There are (5^3)^2 ⋅ 5^0 hens in a bird enclosure. What is the total number of hens in the enclosure?

OpenStudy (horsegal244):

@seascorpion1

OpenStudy (seascorpion1):

Yup

OpenStudy (seascorpion1):

a number to the power 0 equals 1

OpenStudy (horsegal244):

ok

OpenStudy (seascorpion1):

so 5^0=1

OpenStudy (horsegal244):

ok

OpenStudy (horsegal244):

5^3 is 125

OpenStudy (seascorpion1):

so (5^3)^2=125^2

OpenStudy (seascorpion1):

Can you solve it now?

OpenStudy (horsegal244):

(125^2) • 1 correct

OpenStudy (seascorpion1):

Yes

OpenStudy (horsegal244):

so 250

OpenStudy (seascorpion1):

no, 125*125

OpenStudy (horsegal244):

oh

OpenStudy (seascorpion1):

You done?

OpenStudy (horsegal244):

15,625

OpenStudy (seascorpion1):

Yes

OpenStudy (horsegal244):

ok

OpenStudy (horsegal244):

these are the answers 0 5^5 5^6 5^30

OpenStudy (seascorpion1):

\[(a^b)^c=a^{bc}\]

OpenStudy (horsegal244):

?

OpenStudy (seascorpion1):

(5^3)^2=5^(3*2)=5^6

OpenStudy (horsegal244):

ok

OpenStudy (seascorpion1):

so the answer is 5^6*1=5^6 (the third choice)